a bucket of chicken wings for $12 each and a chicken dinner for $7 each, 28 were sold and $281 dollars was collected. how many buckets of chicken wings and how many dinners were sold?
Answer by stanbon(75887)

a bucket of chicken wings for $12 each and a chicken dinner for $7 each, 28 were sold and $281 dollars was collected. how many buckets of chicken wings and how many dinners were sold?
Equations:
Quantity Eq: w + c = 28
Value Eq:: 12w+ 7c = 281
----
Multiply thru the Quantity Eq. by 7 to get:
7w + 7c = 7*28
---
subtract that from the Value Eq. and solve for "w":
5w = 85
w = 17 (# of wings buckets sold)
---
Since w+c = 28, c = 11 (# of dinner buckets sold)
------------------------
Cheers,
Stan H.